OpenStudy (anonymous):

Find the solution of the differential equation that satisfies the given initial condition. dP/dt=4 square root (Pt), P(1)=2

OpenStudy (jamesj):

Is your equation this: \[ \frac{dP}{dt} = 4\sqrt{P(t)} \] ?

OpenStudy (anonymous):

sqrt(P(t))-2*t-_C1 = 0

OpenStudy (jamesj):

We'd never leave a function in that form ... and you should also show the derivation ;-)

OpenStudy (anonymous):

well plug in initial condition and stuff

OpenStudy (jamesj):

For the record then: write P for P(t). Then the equation we have is \[ \frac{dP}{dt} = 4\sqrt{P} \] Separating variables, \[ \int \frac{dP}{\sqrt{P}} = \int 4 \ dt \] Now integrating we have \[ \frac{1}{2} \sqrt{P} = 4t + C \] or \[ P(t) = (8t + 2C)^2 \] You can now apply the initial condition \( P(1) = 2 \) to simplify this further by solving for \( C \).

OpenStudy (jamesj):

**correction, upon integrating we have \[ 2 \sqrt{P} = 4t + C \] and thus \[ \sqrt{P} = 2t + C/2 \] \[ P(t) = (2t + C/2)^2 \]